### v2

@@ 3.2.1 Assistive Robot
The integration of EmAI into healthcare has led to the development of assistive robots that enhance daily care and support of patients. This section focuses on three key applications, including social assistance, daily living assistance, and mobility assistance.
@@
Figure 7: An EmAI system for socially assistive robotics [590]. It enhances children’s social skills through adaptive interaction. The system assesses and responds to children’s emotional cues and social behaviors by combining high-level planning and low-level actuation.
@@
Studies have indicated that children with ASD often respond positively to robot-assisted therapy, showing increased engagement and reduced anxiety during sessions [591, 592, 593]. Beyond psychological disorders, EmAI systems like ZORA robots [594, 595, 596] have assisted children with severe physical disabilities in achieving therapeutic and educational goals, as well as helping patients with dementia optimize communication. We present a pipeline for an EmAI system that supports children with social disorders, as shown in Figure 7.
@@
Daily helper. For patients facing challenges in independent living, advancements in EmAI have led to the development of various assistive robots designed to support essential daily activities. These robots assist with specific tasks such as eating, dressing, personal hygiene, and medication management. For instance, feeder robots have been developed to help severely disabled patients with self-feeding [597, 598]. These systems integrate infrared sensors to achieve precise spoon control based on the user’s body size and head position, automatically adjusting as needed [599, 600, 601]. This adaptability allows users to independently select food, enhancing their dining experience and overall quality of life. Additionally, robots like ARI [583, 584] are designed to assist in rehabilitation programs. These robots can demonstrate exercises, offer real-time guidance, and provide encouragement, fostering greater patient engagement and adherence to therapeutic regimens [602]. This holistic approach supports not only physical recovery but also the emotional and social well-being of patients, making these assistive technologies invaluable tools in both rehabilitation and independent living.
@@
Locomotion aide. EmAI has significantly advanced mobility and rehabilitation assistance for individuals with physical impairments, offering innovative solutions to restore functionality and support individuals during recovery [603]. For instance, exoskeletons, such as those developed by Ekso Bionics, empower patients with spinal cord injuries to stand and walk, promoting rehabilitation and improving mobility [604]. Similarly, ReWalk [605] enables individuals with paralysis to walk and climb stairs, effectively compensating for physical impairments and delivering positive outcomes. AI-powered wheelchairs demonstrate how EmAI enhances mobility. They use AI algorithms to predict movements and assist users with severe motor or cognitive impairments [606, 607]. These systems analyze past movements, detect obstacles, and enable safe, real-time navigation, helping users intuitively maneuver through complex environments [608, 609]. Furthermore, EmAI-based wheelchairs are noted for their cost-effectiveness, achieved by optimizing control parameters to minimize energy consumption and extend battery life [610]. For mobile balance-assistive robots, they are being developed to help individuals maintain balance and prevent falls during daily activities [611]. Some robots, such as SoloWalk [612], provide weight support as patients walk, further aiding the rehabilitation process. Brain-computer interface-controlled robots, including wheelchair-mounted robotic arms [613] and robotic prosthetics, also play a vital role in enhancing functionality for individuals with paralysis or limb loss, enabling them to perform manipulation tasks and regain independence.
@@ 3.2.2 Companion Robot
Emotional companion. Companion robots for emotional support (mental health) typically operate through two main pathways: virtually EmAI agents [627] and advanced AI-robotics innovations used clinically [628]. Virtually embodied agents, such as chatbots or conversational systems, provide accessible emotional support by engaging users in real-time, empathetic conversations that help relieve loneliness, anxiety, or stress [92, 111]. These virtual companions are easily accessible and offer immediate, low-barrier interaction for those seeking support. In clinical settings, AI-integrated physical robots take mental health support a step further by interacting with users face-to-face, recognizing physical and emotional cues, and adapting responses to offer personalized comfort, social engagement, and therapeutic activities [629]. Together, these virtual and physical AI advancements form a well-rounded support system, meeting both immediate conversational needs and offering in-depth emotional support in clinical environments.
@@
Beyond virtual therapeutic applications, clinicians and researchers are working to bring AI-robotics innovations directly into clinical settings. For instance, intelligent animal-like robots such as Paro, a plush harp seal, are increasingly used to assist patients with dementia [630, 631]. Alongside Paro, the larger eBear is part of a class of “companion bots” designed to serve as at-home health aides, responding to speech and movement with interactive “dialogue” and offering companionship. These robots aim to help elderly, isolated, or depressed individuals through comforting interactions. Several studies have examined the role of such robots in reducing stress, loneliness, and agitation and in improving mood and social connections [632].
@@
Kid health guardian. AI-driven robots and wearable devices aid in diagnosing developmental disorders, monitoring vital signs, and engaging children in interactive exercises [633, 634, 621]. By utilizing AI in physically interactive forms, embodied systems can adjust to a child’s responses in real-time, offering adaptive support that traditional methods may lack. In therapeutic contexts, social robots assist children with autism by facilitating social engagement and enhancing communication skills through controlled, repetitive interactions [615]. Additionally, AI-empowered tools for rehabilitation and physical therapy provide targeted exercises with engaging feedback to support children with motor skill challenges [635]. Overall, EmAI enhances healthcare delivery by offering personalized, interactive, and effective treatments, supporting both physical and cognitive aspects of pediatric care.
@@
AI robots present valuable opportunities for engaging children with ASD, offering a unique approach to social skill development [620]. Studies show that children with autism often respond well to robots, even when they struggle to interact with others [615]. The Kaspar robot, for instance, has shown potential for integration into educational and therapeutic settings [618] to help improve social skills [619]. Early studies show that children with ASDs often respond more positively to robot companions than human therapists, demonstrating increased social behaviors and improved spontaneous language during sessions [616]. The development of social robots also offers a promising approach to supporting children with diabetes [617]. The robot not only assists with health management tasks, such as reminding users to check glucose levels and guiding insulin administration, but also provides emotional support by engaging in friendly, human-like interactions. Designed with input from children and clinicians, the robot fosters social engagement, helping to reduce anxiety around diabetes management and improve adherence to health routines. Early findings suggest positive responses from children, indicating that the robot effectively combines practical health assistance with companionship.
@@
Elderly health caregiver. With aging, the elderly face numerous challenges, including declining physical health and increased vulnerability to chronic diseases. Conditions such as cardiovascular disease, diabetes, and arthritis not only diminish their quality of life but also impose considerable stress on healthcare systems and caregivers. These persistent health issues frequently result in reduced mobility, social isolation, and a dependence on others for daily activities, further exacerbating their vulnerability. In light of these challenges, it is crucial to explore innovative solutions. EmAI-based elderly care robots, with advanced long-term learning capabilities, can adapt to an individual’s health trajectory, habits, and preferences over time. This long-term learning capability may involve the integration of RL for personalized care optimization or meta-learning for rapid adaptation, as well as the application of disease-specific predictive modeling to address the evolving nature of elderly users’ needs.
@@
EmAI has been shown to offer practical and emotional support to enhance the outcomes of elderly patients. Robotics with EmAI capabilities assist elderly individuals by improving mobility, conducting physiological assessments, and monitoring vital signs, which supports the completion of daily activities, rehabilitation exercises, and independent navigation [634]. Studies show that elderly patients feel more secure and comfortable with companion robots, as they can be continuously monitored without intruding on their personal space or autonomy [636]. By combining EmAI monitoring systems, elderly companion robots provide a reliable, non-intrusive way to support patient health, making them a valuable asset in home-based care.
@@
Additionally, these AI-enabled robots enhance the quality of care for dementia patients, providing cognitive engagement and symptom management through structured activities and consistent interactions that traditional care methods may not offer. The majority of elderly individuals with dementia who engaged with social robots Sophie and Jack experienced highly positive interactions [626]. These robots are equipped with human-like communication abilities, emotional expressions, and gestures, and they can play songs, games, and tell stories. They have been shown to provide sensory enrichment and foster social engagement for dementia patients. For instance, many participants were motivated to play bingo with Jack and actively joined other group activities, highlighting the robots’ impact on promoting social and cognitive engagement [625].
@@ 3.3.1 Rescue Robot
EmAI-powered robots are increasingly developed to serve as human assistants or replacements in various high-risk environments. In rescue scenarios, their potential has been widely explored due to their ability to operate safely in dangerous environments and urgent conditions. In the aftermath of natural disasters or life-threatening emergencies, search and rescue robots aim to provide rapid aid and emergency health services to victims. Thanks to the advancement of recent AI approaches including MLLMs and VLMs, EmAI-equipped robots can efficiently locate and assist survivors while supporting rescue operations [642, 643, 644, 645] in disaster-prone areas, such as earthquake zones. A representative workflow of EmAI-based rescue robots is illustrated in the Figure 8.
@@
Figure 8: An EmAI system in rescue robotics [646]: optimizing search and rescue operations. This system integrates sensory data with strategic processing to quickly identify and prioritize victims, effectively navigating through complex environments to facilitate timely and efficient rescue missions.
@@
Life scout. Technological advances have led to sophisticated survivor detection systems, which can quickly identify and locate survivors in emergencies using integrated sensors and drones [647, 648]. These systems leverage thermal and RGB images, and AI to analyze data in real-time, efficiently pinpointing living individuals [649, 650]. Recent developments in AI and robotics have enabled systems to analyze vast amounts of data quickly, making real-time survivor detection more accurate and reliable [651]. By integrating visual and audio signals through a wireless sensor network, established EmAI systems [652, 649] are able to detect survivors’ locations, transmit their position info to a centralized cloud server, and assess environmental risk levels to aid in rescue task planning.
@@
Agile walker. The life-search ability of rescue robots has been enhanced by AI-based mobility systems that are capable of operating in both hazardous and confined environments. In such environments, high mobility is essential for navigating rough terrain, clearing obstacles, and streaming live video from affected areas [653, 654]. Current advancements aim to enable autonomous navigation by equipping these robots with AI-based obstacle detection and avoidance capabilities [655, 656, 657]. Legged robots, particularly quadrupeds equipped with advanced EmAI systems, have demonstrated the ability to navigate rough surfaces, debris, and areas with variable friction, making them highly adaptable in complex terrains [658]. Recent RL-augmented frameworks further enhance robots in adaptive blind locomotion, high-speed movement, dynamic response to obstacles, and adaptability improvement under uncertain conditions [659, 660].
@@
Rescue carrier. At the current stage, most rescue robots remain limited to data retrieval and collection tasks [661, 662, 663, 664]. In the next phase of development, it is crucial to integrate multi-degree-of-freedom manipulators based on advanced EmAI systems, such as VLA models, to further expand the functionality of rescue robots. This includes capabilities like holding, grasping, and transferring objects, as well as providing assistance to victims with medical equipment. Equipped with these medical support tools, rescue robots can now offer direct on-site assistance to victims [665].
@@
Delivery robots are now widely used in healthcare settings, performing in-hospital delivery tasks that were previously handled by humans, thus overcoming limitations related to time constraints and availability. With EmAI systems, they are capable of delivering essential items—such as samples, meals, medications, and medical supplies [666, 667, 668] more efficiently, supporting healthcare facilities in streamlining operations and alleviating staff burdens [669]. These robots can independently navigate complex environments, adapt to dynamic conditions, and carry out functions like obstacle avoidance [670], speech interaction [671], and precise patient facial recognition [672]. During the COVID-19 pandemic, robots have also been deployed for tasks like the delivery of medications and food [673, 674], thereby minimizing human contact and lowering the risk of virus transmission.
@@ 3.4 Biomedical Research
Figure 9: An EmAI system in laboratory automation [683]: enhancing experimental precision and efficiency. This setup utilizes advanced planning algorithms for experiment design and resource management, paired with robotic actuators for precise manipulation of reagents and experimental conditions, ensuring compliance with safety standards.
Automated lab technician. Experiments are critical parts in biomedical research. EmAI-driven robots automate repetitive tasks, increasing throughput and allowing researchers to focus on more complex analysis tasks. Scientists have created robot chemists capable of autonomously conducting complex chemical reactions and analyses [683, 684]. These robotic systems apply EmAI in laboratory automation, enabling them to design experiments [685, 686, 687], execute procedures [688, 689], and interpret results [690, 691, 692] without human intervention. For instance, a platform leverages AI-based synthesis planning, trained on millions of reactions, to propose synthetic routes, which are then fine-tuned by expert chemists and executed on a robotic flow system [693]. This setup allows for scalable, reproducible synthesis of a variety of compounds, such as pharmaceuticals. Another AI-driven robotic chemist also demonstrated its ability to autonomously perform tasks ranging from synthetic planning to executing experiments in batch reactors [694]. In the field of materials science for photo-catalytic hydrogen production, the robot operates autonomously in a laboratory environment, using Bayesian optimization to navigate a complex experimental space with ten variables, performing 688 experiments over eight days [683]. In summary, EmAI can combine precision handling with autonomous decision-making, allowing it to run experiments continuously without human intervention. These experiment setups minimize human bias, accelerate discovery and production, and enable exploration of large experimental spaces.
@@
AI-based knowledge retriever. As EmAI continues to evolve, a growing area of focus is the development of systems that emulate the analytical capabilities of human scientists. Advanced frameworks like ChemCrow [712] and CALMS [713] exemplify this trend, utilizing LLMs such as GPT-4, multimodal generative models, and retrieval approaches like GraphRAG [714] to integrate scientific knowledge, biological principles, and theoretical frameworks [715, 716, 717]. These systems go beyond traditional data processing. They engage in critical thinking by questioning assumptions, evaluating evidence, and validating conclusions. In addition, researchers also integrated large amounts of data from publications in PubMed [718, 719], Wikipedia [720] and other open resources to pretrain generative models. These generative models can function as part of a composite system, integrating EmAI world models, experimental platforms, and human expertise to tackle complex problems that require interdisciplinary approaches [27]. In summary, the development of AI knowledge retriever marks a promising future direction in which EmAI could play an even more integral role in advancing knowledge and innovation across diverse fields.

### v3

@@ 3.2.1 Assistive Robot
The integration of EmAI into healthcare has led to the development of assistive robots that enhance daily care and support for patients. This section focuses on three key applications, including social assistance, daily living assistance, and mobility assistance.
@@
Figure 7: An EmAI system for socially assistive robotics. It enhances children’s social skills through adaptive interaction. The system assesses and responds to children’s emotional cues and social behaviors by combining high-level planning and low-level actuation. Source: Reproduced from [590]. Copyright @2025 ETHealthworld.com.
@@
Studies have indicated that children with ASD often respond positively to robot-assisted therapy, showing increased engagement and reduced anxiety during sessions [591, 592, 593]. Beyond psychological disorders, EmAI systems like ZORA robots [594, 595, 596] have also helped children with severe physical disabilities achieve therapeutic and educational goals, as well as helping patients with dementia improve their communication. We present an EmAI system pipeline that supports children with social disorders, as shown in Figure 7.
@@
Daily helper. For patients facing challenges in independent living, advancements in EmAI have led to the development of various assistive robots designed to support essential daily activities. These robots assist with specific tasks such as eating, dressing, personal hygiene, and medication management. For instance, feeder robots have been developed to help severely disabled patients with self-feeding [597, 598]. These systems integrate infrared sensors to achieve precise spoon control based on the user’s body size and head position, automatically adjusting as needed [599, 600, 601]. This adaptability allows users to independently select food, enhancing their dining experience and overall quality of life. Additionally, robots like ARI [583, 584] are designed to assist in rehabilitation programs. These robots can demonstrate exercises, offer real-time guidance, and provide encouragement, fostering greater patient engagement and adherence to therapeutic regimens [602]. By supporting both physical recovery and emotional well-being, these assistive technologies are becoming essential in helping patients regain independence and improve their quality of life.
@@
Locomotion aide. EmAI has significantly advanced mobility and rehabilitation assistance for individuals with physical impairments, offering innovative solutions to restore functionality and support individuals during recovery [603]. For instance, exoskeletons, such as those developed by Ekso Bionics, empower patients with spinal cord injuries to stand and walk, promoting rehabilitation and improving mobility [604]. Similarly, ReWalk [605] enables individuals with paralysis to walk and climb stairs, effectively compensating for physical impairments and delivering positive outcomes. AI-powered wheelchairs demonstrate how EmAI enhances mobility. They use AI algorithms to predict movements and assist users with severe motor or cognitive impairments [606, 607]. These systems analyze past movements, detect obstacles, and enable safe, real-time navigation, helping users intuitively maneuver through complex environments [608, 609]. Additionally, EmAI-based wheelchairs are noted for their cost-effectiveness, achieved by optimizing control parameters to minimize energy consumption and extend battery life [610]. For mobile balance-assistive robots, these systems are being developed to help individuals maintain balance and prevent falls during daily activities [611]. Some robots, such as SoloWalk [612], provide weight support as patients walk, further aiding the rehabilitation process. Brain-computer interface-controlled robots, including wheelchair-mounted robotic arms [613] and robotic prosthetics, also play a vital role in enhancing functionality for individuals with paralysis or limb loss, enabling them to perform manipulation tasks and regain independence.
@@ 3.2.2 Companion Robot
Emotional companion. Companion robots for emotional support typically operate through two main pathways: virtually EmAI agents [627] and advanced AI-robotics innovations used clinically [628]. Virtually embodied agents, such as chatbots or conversational systems, provide accessible emotional support by engaging users in real-time, empathetic conversations that help relieve loneliness, anxiety, or stress [92, 111]. These virtual companions are easily accessible and offer immediate, low-barrier interaction for those seeking support. In clinical settings, AI-integrated physical robots take mental health support a step further by interacting with users face-to-face, recognizing physical and emotional cues, and adapting responses to offer personalized comfort, social engagement, and therapeutic activities [629]. Together, these virtual and physical AI advancements form a well-rounded support system, meeting both immediate conversational needs and offering in-depth emotional support in clinical environments.
@@
Beyond virtual therapeutic applications, clinicians and researchers are working to bring AI-robotics innovations directly into clinical settings. For instance, intelligent animal-like robots, such as Paro, a plush harp seal, are increasingly used to assist patients with dementia [630, 631]. Alongside Paro, the larger eBear is part of a class of “companion bots” designed to serve as at-home health aides, responding to speech and movement with interactive “dialogue” and offering companionship. These robots aim to help elderly, isolated, or depressed individuals through comforting interactions. Several studies have examined the role of such robots in reducing stress, loneliness, and agitation and in improving mood and social connections [632].
@@
Kid health guardian. AI-driven robots and wearable devices play a key role in diagnosing developmental disorders, monitoring vital signs, and engaging children in interactive exercises [633, 634, 621]. By incorporating AI into interactive, physical forms, these embodied systems can adjust to a child’s responses in real-time, offering adaptive support that traditional methods may lack. In therapeutic contexts, social robots help children with autism by promoting social engagement and enhancing communication skills through controlled, repetitive interactions [615]. Additionally, AI-based tools for rehabilitation and physical therapy provide targeted exercises with engaging feedback to support children with motor skill challenges [635]. Overall, EmAI is revolutionizing pediatric healthcare by delivering personalized, interactive, and effective treatments that address both the physical and cognitive needs of children.
@@
AI robots present valuable opportunities for engaging children with ASD, offering a unique approach to social skill development [620]. Studies show that children with autism often respond well to robots, even when they struggle to interact with others [615]. The Kaspar robot, for instance, has shown potential for integration into educational and therapeutic settings [618] to help improve social skills [619]. Early studies suggest that children with ASDs often respond more positively to robot companions than human therapists, demonstrating increased social behaviors and improved spontaneous language during sessions [616]. Furthermore, the development of social robots also offers a promising approach to supporting children with diabetes [617]. Such robots not only assist with health management tasks, such as reminding users to check glucose levels and guiding insulin administration, but also provide emotional support by engaging in friendly, human-like interactions. Designed with input from children and clinicians, the robot fosters social engagement, reducing anxiety around diabetes management and improving adherence to health routines. Early findings are encouraging, showing that these robots effectively combine practical health assistance with the benefits of companionship.
@@
Elderly health caregiver. With aging, the elderly face numerous challenges, including declining physical health and increased vulnerability to chronic diseases. Conditions such as cardiovascular disease, diabetes, and arthritis not only diminish their quality of life but also impose considerable stress on healthcare systems and caregivers. These persistent health issues frequently result in reduced mobility, social isolation, and a dependence on others for daily activities, further exacerbating their vulnerability. In light of these challenges, it is crucial to explore innovative solutions. EmAI-based elderly care robots, with advanced long-term learning capabilities, can adapt to an individual’s health trajectory, habits, and preferences over time. This long-term learning capability may include integrating RL for personalized care optimization or meta-learning for rapid adaptation, as well as the application of disease-specific predictive modeling to address the evolving nature of elderly users’ needs.
@@
EmAI has been shown to offer practical and emotional support to enhance the outcomes of elderly patients. Robotics with EmAI capabilities assist elderly individuals by improving mobility, conducting physiological assessments, and monitoring vital signs, which helps with daily activities, rehabilitation exercises, and independent navigation [634]. Studies show that elderly patients feel more secure and comfortable with companion robots, as they can be continuously monitored without intruding on their personal space or autonomy [636]. By combining EmAI monitoring systems, elderly companion robots provide a reliable, non-intrusive way to support patient health, making them a valuable asset in home-based care.
@@
Additionally, these AI-enabled robots enhance the quality of care for dementia patients, providing cognitive engagement and symptom management through structured activities and consistent interactions that traditional care methods may not be able to offer. The majority of elderly individuals with dementia who engaged with social robots Sophie and Jack experienced highly positive interactions [626]. These robots are equipped with human-like communication abilities, emotional expressions, and gestures, and they can play songs, games, and tell stories. They have been shown to provide sensory enrichment and foster social engagement for dementia patients. For instance, many participants were motivated to play bingo with Jack and actively joined other group activities, highlighting the robots’ impact on promoting social and cognitive engagement [625].
@@ 3.3.1 Rescue Robot
EmAI-based robots are increasingly being developed to serve as human assistants or replacements in various high-risk environments. In rescue scenarios, their potential has been widely explored due to their ability to operate safely in dangerous environments and urgent conditions. In the aftermath of natural disasters or life-threatening emergencies, search and rescue robots aim to provide rapid aid and emergency health services to victims. Thanks to the advancement of recent AI approaches including MLLMs and VLMs, EmAI-equipped robots can efficiently locate and assist survivors while supporting rescue operations [642, 643, 644, 645] in disaster-prone areas, such as earthquake zones. A representative workflow of EmAI-based rescue robots is illustrated in Figure 8.
@@
Figure 8: An EmAI system in rescue robotics: optimizing search and rescue operations. This system integrates sensory data with strategic processing to quickly identify and prioritize victims, effectively navigating through complex environments to facilitate timely and efficient rescue missions. Source: Reproduced from [646]. Copyright ©2024 Boston Dynamics.
@@
Life scout. Technological advances have produced sophisticated survivor detection systems, which can quickly identify and locate survivors in emergencies using integrated sensors and drones [647, 648]. These systems leverage thermal and RGB imaging, along with AI for real-time data analysis, efficiently pinpointing survivors [649, 650]. Recent developments in AI and robotics have enabled systems to analyze vast amounts of data quickly, making real-time survivor detection more accurate and reliable [651]. By integrating both visual and audio signals through a wireless sensor network, established EmAI systems [652, 649] are able to detect survivors’ locations, transmit their position information to a centralized cloud server, and assess environmental risk levels to support rescue planning.
@@
Agile walker. The life-search ability of rescue robots has been enhanced by AI-based mobility systems that are capable of operating in both hazardous and confined environments. In such environments, high mobility is essential for navigating rough terrain, clearing obstacles, and streaming live video from affected areas [653, 654]. Current advancements aim to enable autonomous navigation by equipping these robots with AI-based obstacle detection and avoidance capabilities [655, 656, 657]. Legged robots, particularly quadrupeds equipped with advanced EmAI systems, have demonstrated the ability to navigate rough surfaces, debris, and areas with variable friction, making them highly adaptable in complex terrains [658]. Furthermore, recent RL-augmented frameworks further enhance robots in adaptive blind locomotion, high-speed movement, dynamic response to obstacles, and adaptability improvement under uncertain conditions [659, 660].
@@
Rescue carrier. At the current stage, most rescue robots remain limited to data retrieval and collection tasks [661, 662, 663, 664]. In the next phase, it is crucial to integrate multi-degree-of-freedom manipulators based on advanced EmAI systems, such as VLA models, to further expand the functionality of rescue robots. These enhancements would enable them to hold, grasp, and transfer objects, as well as assist victims with medical equipment. With these medical support tools, rescue robots could offer direct on-site assistance to victims [665].
@@
Delivery robots are now widely used in healthcare settings, performing in-hospital delivery tasks that were previously handled by humans, thus overcoming limitations related to time constraints and availability. With EmAI systems, they are capable of delivering essential items—such as samples, meals, medications, and medical supplies [666, 667, 668] more efficiently, supporting healthcare facilities in streamlining operations and alleviating staff burdens [669]. These robots can independently navigate complex environments, adapt to dynamic conditions, and carry out functions like obstacle avoidance [670], speech interaction [671], and precise patient facial recognition [672]. During the COVID-19 pandemic, robots have also been deployed for tasks like the delivery of medications and food [673, 674], thereby minimizing human contact and reducing the risk of virus transmission.
@@ 3.4 Biomedical Research
Figure 9: An EmAI system in laboratory automation: enhancing experimental precision and efficiency. This setup utilizes advanced planning algorithms for experiment design and resource management, paired with robotic actuators for precise manipulation of reagents and experimental conditions, ensuring compliance with safety standards. Source: Reproduced from [683] with permission from Springer Nature.
Automated lab technician. Experiments are critical components in biomedical research. EmAI-driven robots automate repetitive tasks, increasing throughput and allowing researchers to focus on more complex analysis tasks. Scientists have created robot chemists capable of autonomously conducting complex chemical reactions and analyses [683, 684]. These robotic systems apply EmAI in laboratory automation, enabling them to design experiments [685, 686, 687], execute procedures [688, 689], and interpret results [690, 691, 692] without human intervention. For instance, a platform leverages AI-based synthesis planning, trained on millions of reactions, to propose synthetic routes, which are then fine-tuned by expert chemists and executed on a robotic flow system [693]. This setup allows for scalable, reproducible synthesis of a variety of compounds, such as pharmaceuticals. Another AI-driven robotic chemist also demonstrated its ability to autonomously perform tasks ranging from synthetic planning to executing experiments in batch reactors [694]. In the field of materials science for photo-catalytic hydrogen production, the robot operates autonomously in a laboratory environment, using Bayesian optimization to navigate a complex experimental space with ten variables, performing 688 experiments over eight days [683]. In summary, EmAI can combine precision handling with autonomous decision-making, allowing it to run experiments continuously without human intervention. These experiment setups minimize human bias, accelerate discovery and production, and enable exploration of large experimental spaces.
@@
AI-based knowledge retriever. As EmAI continues to evolve, an increasingly important focus is on developing systems that mimic the analytical capabilities of human scientists. Advanced frameworks like ChemCrow [712] and CALMS [713] exemplify this trend, utilizing LLMs such as GPT-4, multimodal generative models, and retrieval approaches like GraphRAG [714] to integrate scientific knowledge, biological principles, and theoretical frameworks [715, 716, 717]. These systems go beyond traditional data processing. They engage in critical thinking by questioning assumptions, evaluating evidence, and validating conclusions. In addition, researchers have integrated large amounts of data from publications in PubMed [718, 719], Wikipedia [720] and other open resources to pretrain generative models. These generative models can function as part of a composite system, integrating EmAI world models, experimental platforms, and human expertise to tackle complex problems that require interdisciplinary approaches [74]. In summary, the development of AI knowledge retriever marks a promising future direction in which EmAI could play an even more integral role in advancing knowledge and innovation across diverse fields.